OpenStudy (anonymous):

HELP!!! PLEASE! Samantha received 6 dolls for her birthday. She plans to take two of them with her to grandma’s house. How many combinations of 2 dolls from the group of 6 could she take with her?

OpenStudy (anonymous):

Do you know nCr? I think that's what you're supposed to use.

OpenStudy (anonymous):

I do not know what that is

OpenStudy (anonymous):

The binomial coefficient? Pascal's triangle?

OpenStudy (shamil98):

x + y + z + n + k + h let's use variables how many different groups can you get from this? xy yz zn nk kh xz yn zk nh xn yk zh xk yh xk

OpenStudy (shamil98):

5 + 4 + 3 + 2 + 1 = 15

OpenStudy (wolf1728):

The formula for selecting k objects from n objects is combinations = n! /(k! * (n-k)! combinations = 6*5*4*3*2*1 / 2! * (6-2)! combinations = 6*5*4*3*2*1 / 2 * 4 * 3 * 2 * 1 =6 * 5 / 2 = 15

OpenStudy (shamil98):

Yes, you can use that as well, I just thought a simple algebraic example would be easier. But my method is not that efficient when it comes to larger values of n.

OpenStudy (shamil98):

Which is why one would use the formula wolf put out.

OpenStudy (wolf1728):

That's the reason I showed it. (not to upstage anyone) Also, if the teacher wants an answer using a formula there it is. :-)